# Selecting informative conformal prediction sets with false coverage rate control

### v1

@@ 6 Application to directional FDR control
Consider for this section that we have $(X_{i},Z_{i})_{i\in[n+m]}$ with real-valued outcomes $Z_{i}\in\mathbb{R}$, and we aim at deciding if $Z_{n+i}\in[a,b]$, $Z_{n+i}<a$ or $Z_{n+i}>b$ (without producing prediction intervals for the $Z_{n+i}$), for two benchmark values $a\leq b$. More formally, we want to build a selection $\mathcal{S}\subset[m]$ and a (point-wise) decision $\widehat{Y}\in\{1,2,3\}$ from the observed samples such that
@@
We consider =$\mathcal{I}=\{C\subset[K]\>:\>y_{0}\notin C,|C|\leq 1\}$ for $y_{0}=2$ (see Example 2.3) in the classification setting based on the sample $(X_{j},Y_{j})$’s ($K=3$), and we note that the FCR coincides with the directional FDR in that case. Hence, the result comes directly from Theorem 3.1 (note that InfoSP is post-processed here, see § 5.1). ∎
Remark 6.1.
@@
For $b=a$ and continuous outcomes (with, say, $\tilde{p}^{(2)}_{i}=0$), there are almost surely only two classes, $Z_{i}=1$ corresponding to $Y_{i}<a$ and $Z_{i}=3$ corresponding to $Y_{i}>a$. The procedure is thus InfoSP for non-trivial classification, with class-calibrated $p$-values for $K=2$. This procedure coincides with the directional FDR procedure in Guo and Romano, (2015) applied to conformal $p$-values. The proof of validity in Guo and Romano, (2015) assumes that the test statistics for the $m$ hypotheses are independent. Interestingly, with the dependence induced by the conformal $p$-values, the same procedure is still valid, as formalized in Corollary 6.1.
7 Informative prediction sets for 3 classes of animals
In this section, we illustrate the performance of our methods on real data. We use the image dataset CIFAR-10 (https://www.cs.toronto.edu/~kriz/cifar.html), which consists of 60000 32x32 colour images in 10 classes, with 6000 images per class. We restrict the analysis to 3 classes: birds, cats and dogs.
We consider 4 scenarios: a) non-trivial classification task in a i.i.d. setting, b) non-trivial classification task in a class-conditional setting, c) non-null classification task in a i.i.d. setting, and d) non-null classification task in a class-conditional setting (in c) and d), the null class is the bird class). Specifically, while the classes are in equal proportions by default, we introduce a label shift between the calibration sample and the test sample for the class-conditional settings, in b) by modifying the classes proportions of the calibration sample to 20% for the bird class versus 40% for the cat/dog class and in d) by modifying the classes proportions of the test sample to 50% for the bird class versus 25% for the cat/dog class. In each experiment, the test size is $m=5000$ and the calibration size is $n=5000$.
In the i.i.d. settings (scenarios a) and c)), we evaluate the procedures InfoSP and InfoSCOP with full-calibrated $p$-values and in the class-conditional settings (scenarios b) and d)) we evaluate InfoSP with class-calibrated $p$-values, each being also compared with classical conformal prediction (denoted by CC). For all procedures the non-conformity score is $S_{y}(x)=1-\pi_{y}(x)$ where $\pi_{y}(x)$ is an estimator of the probability that the class of $x$ is $y$ and is learned using a convolutional neural network (CNN) with 2 convolutional layers, one pooling layer, and 3 fully-connected layers, trained for 20 epochs with a learning rate of 0.01 on a sample size of 5000. To assess the power of the methods, in addition to the resolution-adjusted power, we plot the selection rate (SR) of the procedures, defined as the average proportion of informative prediction sets returned, and the average length of the informative prediction sets.
In each setting, the FDR and each power metric for the methods are evaluated by using 100 runs and the results are reported in Figure 8 for $\alpha=0.1$. The conclusions are qualitatively similar to the experiments of § 5: in all settings, classical conformal prediction yields a FCR that severely exceeds the marginal level with an inflation of about 50%. By contrast, our procedures InfoSP and InfoSCOP control the FCR at the target nominal level, both without and with label shift. In terms of power, concerning the i.i.d. settings, pre-processing is not useful for non-trivial classification as expected and the performances of InfoSP and InfoSCOP are similar with a FCR close to the nominal level for both in that case. Conversely, for the non-null classification task, InfoSP is conservative while InfoSCOP is more powerful and displays a FCR close to $\alpha$. When there is label shift, in the case of non-trivial classification InfoSP displays a FCR close to $\alpha$. In the non-null classification case, however, the label shift increases the difficulty of the task in the sense that for a fixed selection, under-covering the null class in the test sample results in more false informative sets. Hence, the power is low in that case. Finally, in all scenarios, while our procedures InfoSP and InfoSCOP output a lower number of informative sets compared to CC, which is necessary in order to control the FCR, the average size of the prediction sets that are informative is comparable.
a) Non-trivial classification
b) Non-trivial classification with label shift
c) Non-null classification
d) Non-null classification with label shift
Figure 8: FCR, SR, average length on selection and resolution-adjusted power for the methods, for $\alpha=0.1$.
8 Conclusion and discussion
Our methods are very general, and they are relevant to applications in classification and regression. We showed examples in § 4-§ 7 for informative subsets of interest. We used common scores for the examples, but many other scores from the literature can be used with our suggested procedures InfoSP and InfoSCOP as long as Assumption 1 is satisfied. In addition, transfer learning scores can also be handled by our theory (see Assumption 6 in SM). This is known to greatly improve the conformal inference when there is a domain shift between the learning sample and the calibration+test samples (Courty et al.,, 2017; Gazin et al.,, 2023) and developing specific application cases for the latter is of interest for future investigations.
For the iid model, InfoSCOP improves over InfoSP in all considered examples, except when selecting non-trivial prediction subsets in classification (Example 1.2 item 2), for which we establish that InfoSP almost exhausts the FCR level (see Proposition 5.1for $K=2$). The procedure InfoSCOP splits the calibration sample in order to apply an efficient initial selection step on part of the calibration sample and the test sample. When the calibration sample is small, it may be of interest to examine whether alternative methods (e.g., relevant variants of the Jackknife Barber et al.,, 2021) may be incorporated into our framework in settings where InfoSCOP improves over InfoSP but splitting may be too costly. When the calibration sample is large enough, as in our numerical examples, splitting the calibration sample for the initial selection $S^{(0)}$ worked well. However, there are many possibilities about how to perform the initial selection. The choice is important because it defines the pre-processed $p$-values (24) that can be seen as $p$-values “conditionally on being selected”. Different ways of choosing $S^{(0)}$ have been investigated: trying to rule out all the examples in null class (§ 5.3) or trying to mimic the $\mathrm{BH}(\mathbf{q})$ selection that will be applied at the second stage to reduce the selection effect (§ 4). Finding an optimal way of calibrating $S^{(0)}$ is an interesting avenue for future research.
We provided a class conditional variant of InfoSP, with class-conditional guarantees. We proved that our strategy can be followed in the case where the classes of the calibration and test samples are arbitrary fixed, even when the class proportions in the calibration are very different than in the test. The main point is that working with the class-calibrated $p$-value collection allows to maintain the FCR control in this strong sense. In § B we suggest additionally weighted procedures, that incorporate the estimated class proportions. These procedures are not necessarily more powerful than InfoSP, and further research is needed in order to make recommendations about when to use the weighted procedures, and about weight adaptation to the specifics of the data.
Acknowledgments
The authors acknowledge the grants ANR-21-CE23-0035 (ASCAI) and ANR-23-CE40-0018-01 (BACKUP) of the French National Research Agency ANR, and the Emergence project MARS of Sorbonne Université. The second author was supported in part by Israeli Science Foundation grant no. 2180/20.
Appendix A FCR control for BY-type selection rules
@@
In this section, we present a general approach for FCR control (§ A.1), which relies on the following:
@@
  • a general class of $p$-values, including both iid and class-conditional conformal $p$-values (§ A.3);
  • a general selection rule, including informative selection rules (§ A.4).
The assumption on the selection rule is inspired by Benjamini and Yekutieli, (2005); Benjamini and Bogomolov, (2013).
A.1 General statement
@@
Let $m\geq 1$, $\mathcal{Y}\subset\mathbb{R}$ and consider a family $\mathbf{p}=\{p^{(y)}_{i},i\in[m],y\in\mathcal{Y}\}$ of random variables taking values in $[0,1]$ and a vector $Y=(Y_{n+i},i\in[m])$ taking values in $\mathcal{Y}$. We make use of the notation $\mathbf{p}_{-i}:=(p^{(y)}_{j})_{j\neq i,y\in\mathcal{Y}}$ for all $i\in[m]$.
@@
Second, for any selection rule $\mathcal{S}\subset[m]$, that is, any measurable function of $\mathbf{p}$ valued in the subsets of $[m]$, we introduce the quantity (similar to Benjamini and Yekutieli, (2005); Bao et al., (2024))
$s_{i}^{\min}(\mathbf{p}_{-i})=\min_{z\in[0,1]^{\mathcal{Y}}:i\in\mathcal{S}(z,\mathbf{p}_{-i})}\left|\mathcal{S}(z,\mathbf{p}_{-i})\right|\geq 1,\>\>\>i\in[m],$ (28)
with the following assumption.
@@
For $i\in[m]$, $s_{i}^{\min}(\mathbf{p}_{-i})$ is a coordinate-wise nonincreasing function of $\mathbf{p}_{-i}$.
Theorem A.1.
Let us consider a $p$-value family $\mathbf{p}$ and a label/outcome vector $Y$ satisfying Assumption 4, for a selection rule $\mathcal{S}\subset[m]$ with $s_{i}^{\min}=s_{i}^{\min}(\mathbf{p}_{-i})$ (28) satisfying Assumption 5. Then, the procedure $\mathcal{R}_{\alpha}=\big{(}\mathcal{C}^{\alpha s_{i}^{\min}/m}_{n+i}(\mathbf{p})\big{)}_{i\in\mathcal{S}}$, for which the prediction set is defined as in (4) with level $\alpha s_{i}^{\min}/m$, satisfies $\mathbb{E}(\mathrm{FCP}(\mathcal{R}_{\alpha},Y))\leq\alpha,$ for which the expectation $\mathbb{E}$ is taken w.r.t. the same probability than the one of (27) in Assumption 4 (ii).
The inequality $\mathbb{E}(\mathrm{FCP}(\mathcal{R}_{\alpha},Y))\leq\alpha$ in Theorem A.1 will be used both in the cases where $Y_{n+i}$ is fixed (conditional model) or not (iid model). The proof is provided in § A.2. Note that the considered assumptions make the proof particularly simple. Assumptions 4 and 5 are studied in § A.3 and § A.4, respectively.
A.2 Proof of Theorem A.1
@@
$\mathrm{FCP}(\mathcal{R}_{\alpha},Y)$ $=\frac{\sum_{i\in\mathcal{S}}{\mathbf{1}\{Y_{n+i}\notin\mathcal{C}_{n+i}^{\alpha s_{i}^{\min}/m}(\mathbf{p})\}}}{1\vee|\mathcal{S}|}=\sum_{i\in[m]}\frac{{\mathbf{1}\{i\in\mathcal{S},p_{i}^{(Y_{n+i})}\leq\alpha s^{\min}_{i}/m\}}}{1\vee|\mathcal{S}|}$
$\leq\sum_{i\in[m]}\frac{{\mathbf{1}\{i\in\mathcal{S},p_{i}^{(Y_{n+i})}\leq\alpha s^{\min}_{i}/m\}}}{s^{\min}_{i}}\leq\sum_{i\in[m]}\frac{{\mathbf{1}\{p_{i}^{(Y_{n+i})}\leq\alpha s^{\min}_{i}/m\}}}{s^{\min}_{i}},$
@@
$\mathbb{E}\left(\mathrm{FCP}(\mathcal{R}_{\alpha},Y)\right)$ $\leq\sum_{i\in[m]}\mathbb{E}\left(\frac{{\mathbf{1}\{p_{i}^{(Y_{n+i})}\leq\alpha s^{\min}_{i}/m\}}}{s^{\min}_{i}}\right)$
$=\sum_{i\in[m]}\mathbb{E}\left(\mathbb{E}\left[\frac{{\mathbf{1}\{p_{i}^{(Y_{n+i})}\leq\alpha s^{\min}_{i}(\mathbf{p}_{-i})/m\}}}{s^{\min}_{i}(\mathbf{p}_{-i})}\>\middle|\>W_{i}\right]\right),$
@@
$s^{\min}_{i}(\mathbf{p}_{-i})=s^{\min}_{i}(\Psi_{i}(p^{(Y_{n+i})}_{i},W_{i}))$
is a nonincreasing function of $p^{(Y_{n+i})}_{i}$. By Assumption 4 (ii) and applying Lemma D.1 (conditionally on $W_{i}$), we obtain
$\mathbb{E}\left[\frac{{\mathbf{1}\{p_{i}^{(Y_{n+i})}\leq\alpha s^{\min}_{i}(\mathbf{p}_{-i})/m\}}}{s^{\min}_{i}(\mathbf{p}_{-i})}\>\middle|\>W_{i}\right]\leq\frac{\alpha}{m},$
@@
A.3 Examining Assumption 4
We show here that the $p$-value families we consider satisfy Assumption 4 both in the iid and conditional model. For this, it is interesting to relax Assumption 1 by assuming that the score functions can use the covariates of the calibration+test samples in an exchangeable way, as suggested in Marandon et al., (2024); Gazin et al., (2023). This is useful for instance when the learning sample and the calibration+test samples are not based on the same distribution, so that the scores may be improved by using transfer learning; we refer to Gazin et al., (2023) for more details on this.

### v3

@@ Appendix B Application to directional FDR control
Consider for this section that we have $(X_{i},Z_{i})_{i\in[n+m]}$ with real-valued outcomes $Z_{i}\in\mathbb{R}$, and we aim at excluding $Z_{n+i}\in[a,b]$, as well as at deciding whether $Z_{n+i}<a$ or $Z_{n+i}>b$ (without producing prediction intervals for the $Z_{n+i}$), for two benchmark values $a\leq b$. More formally, we want to build a selection $\mathcal{S}\subset[m]$ and a (point-wise) decision $\widehat{Y}\in\{1,3\}$ from the observed samples such that
@@
We consider =$\mathcal{I}=\{C\subset[K]\>:\>y_{0}\notin C,|C|\leq 1\}$ for $y_{0}=2$ (see Example 2.3) in the classification setting based on the sample $(X_{j},Y_{j})$’s ($K=3$), and we note that the FCR coincides with the directional FDR in that case. Hence, the result comes directly from Theorem 3.1 (note that InfoSP is post-processed here, see § 5). ∎
Remark B.1.
@@
For $b=a$ and continuous outcomes (with, say, $\tilde{p}^{(2)}_{i}=0$), there are almost surely only two classes, $Z_{i}=1$ corresponding to $Y_{i}<a$ and $Z_{i}=3$ corresponding to $Y_{i}>a$. The procedure is thus InfoSP for non-trivial classification, with class-calibrated $p$-values for $K=2$. This procedure coincides with the directional FDR procedure in Guo and Romano, (2015) applied to conformal $p$-values for the parameters $Y_{n+1},\ldots,Y_{n+m}$. The proof of validity in Guo and Romano, (2015) assumes that the test statistics for the $m$ hypotheses are independent. Interestingly, with the dependence induced by the conformal $p$-values, the same procedure is still valid, as formalized in Corollary B.1.
Remark B.2.
If one wants to obtain prediction intervals in addition to the directional FDR control, it turns out that in the setting of Corollary 4.1, not only the FDR control (ii) holds but also the directional FDR control
$\mathrm{FDR}_{{\tiny\mbox{dir}}}(\mathcal{R}):=\sup_{P_{X,Y}}\mathbb{E}_{(X,Y)\sim P_{X,Y}}\left[\frac{\sum_{i\in\mathcal{S}}{\mathbf{1}\{D_{i}=1,Y_{n+i}>a\}}+{\mathbf{1}\{D_{i}=3,Y_{n+i}<b\}}}{1\vee|\mathcal{S}|}\right]\leq\alpha,$
for the procedure $\mathcal{R}=(\mathcal{C}_{n+i})_{i\in\mathcal{S}}$ defined therein with the directional rule $D_{i}={\mathbf{1}\{\mathcal{C}_{n+i}\subset(-\infty,a)\}}+3{\mathbf{1}\{\mathcal{C}_{n+i}\subset(b,+\infty)\}}$ for $i\in\mathcal{S}$. It can be slightly less powerful than the directional FDR controlling procedure of Corollary B.1 (because the latter uses classification scores), but provides additional information.
Appendix C FCR control for concordant selection rules
@@
In this section, we present a general approach for FCR control (§ C.1), which relies on the following:
@@
  • a general class of $p$-values, including both full-calibrated and class-calibrated conformal $p$-values (§ C.3);
  • the class of concordant selection rules (Benjamini and Yekutieli,, 2005; Benjamini and Bogomolov,, 2013), including informative selection rules (§ C.4).
C.1 General statement
@@
Let $m\geq 1$, $\mathcal{Y}\subset\mathbb{R}$ and consider a family $\mathbf{p}=(p^{(y)}_{i},i\in[m],y\in\mathcal{Y})$ of random variables taking values in $[0,1]$ and a vector $Y=(Y_{n+i},i\in[m])$ taking values in $\mathcal{Y}$. We make use of the notation $\mathbf{p}_{-i}:=(p^{(y)}_{j})_{j\neq i,y\in\mathcal{Y}}$ for all $i\in[m]$.
@@
Second, for any selection rule $\mathcal{S}\subset[m]$, that is, any measurable function of $\mathbf{p}$ valued in the subsets of $[m]$, we introduce the following quantity (Benjamini and Yekutieli,, 2005; Bao et al.,, 2024):
$s_{i}^{\min}(\mathbf{p}_{-i})=\min_{z\in\mathcal{A}(\mathbf{p}_{-i})}\left|\mathcal{S}(z,\mathbf{p}_{-i})\right|,\>\>\>i\in[m],$ (26)
where $\mathcal{A}(\mathbf{p}_{-i})=\{z\in[0,1]^{\mathcal{Y}}:i\in\mathcal{S}(z,\mathbf{p}_{-i})\}$ (and by convention $s_{i}^{\min}(\mathbf{p}_{-i})=0$ if $\mathcal{A}(\mathbf{p}_{-i})$ is empty). The following assumption corresponds to the concordance assumption of Benjamini and Bogomolov, (2013) (itself extending the former concordance assumption of Benjamini and Yekutieli,, 2005 to a collection of $p$-value families).
@@
For all $i\in[m]$, $s_{i}^{\min}(\mathbf{p}_{-i})$ is a coordinate-wise nonincreasing function of $\mathbf{p}_{-i}$.
As we show in Section C.4, $\mathcal{I}$-informative selection rules are concordant. Other examples include monotone $p$-value-based thresholding rules. However, top-$k$ type selections are concordant only if we use a tie-breaker in the conformal $p$-value definitions (that is, if we use randomized conformal $p$-values Vovk et al.,, 2005; Bates et al.,, 2023; Jin and Candes,, 2023). More formal details are provided in Section C.4.
Theorem C.1.
Let us consider a $p$-value family $\mathbf{p}$ and a label/outcome vector $Y$ satisfying Assumption 4, for a selection rule $\mathcal{S}\subset[m]$ with $s_{i}^{\min}=s_{i}^{\min}(\mathbf{p}_{-i})$ (26) satisfying Assumption 5. Then, the procedure $\mathcal{R}_{\alpha}=\big{(}\mathcal{C}^{\alpha{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}(1\vee s_{i}^{\min})}/m}_{n+i}(\mathbf{p})\big{)}_{i\in\mathcal{S}}$, for which the prediction set is defined as in (4) with level $\alpha(1\vee s_{i}^{\min})/m$, satisfies $\mathbb{E}(\mathrm{FCP}(\mathcal{R}_{\alpha},Y))\leq\alpha,$ for which the expectation $\mathbb{E}$ is taken w.r.t. the same probability as the one of (25) in Assumption 4 (ii).
The inequality $\mathbb{E}(\mathrm{FCP}(\mathcal{R}_{\alpha},Y))\leq\alpha$ in Theorem C.1 will be used both in the cases where $Y_{n+i}$ is fixed (conditional model) or not (iid model). The proof is provided in § C.2. Note that the considered assumptions make the proof particularly simple. Assumptions 4 and 5 are studied in § C.3 and § C.4, respectively.
C.2 Proof of Theorem C.1
@@
$\mathrm{FCP}(\mathcal{R}_{\alpha},Y)$ $=\frac{\sum_{i\in\mathcal{S}}{\mathbf{1}\{Y_{n+i}\notin\mathcal{C}_{n+i}^{\alpha({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}1\vee s_{i}^{\min}})/m}(\mathbf{p})\}}}{1\vee|\mathcal{S}|}=\sum_{i\in[m]}\frac{{\mathbf{1}\{i\in\mathcal{S},p_{i}^{(Y_{n+i})}\leq\alpha({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}1\vee s^{\min}_{i}})/m\}}}{1\vee|\mathcal{S}|}$
$\leq\sum_{i\in[m]}\frac{{\mathbf{1}\{i\in\mathcal{S},p_{i}^{(Y_{n+i})}\leq\alpha({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}1\vee s^{\min}_{i}})/m\}}}{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}1\vee s^{\min}_{i}}}\leq\sum_{i\in[m]}\frac{{\mathbf{1}\{p_{i}^{(Y_{n+i})}\leq\alpha({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}1\vee s^{\min}_{i}})/m\}}}{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}1\vee s^{\min}_{i}}},$
@@
$\mathbb{E}\left(\mathrm{FCP}(\mathcal{R}_{\alpha},Y)\right)$ $\leq\sum_{i\in[m]}\mathbb{E}\left(\frac{{\mathbf{1}\{p_{i}^{(Y_{n+i})}\leq\alpha({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}1\vee s^{\min}_{i}})/m\}}}{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}1\vee s^{\min}_{i}}}\right)$
$=\sum_{i\in[m]}\mathbb{E}\left(\mathbb{E}\left[\frac{{\mathbf{1}\{p_{i}^{(Y_{n+i})}\leq\alpha({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}1\vee s^{\min}_{i}}(\mathbf{p}_{-i}))/m\}}}{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}1\vee s^{\min}_{i}}(\mathbf{p}_{-i})}\>\middle|\>W_{i}\right]\right),$
@@
${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}1\vee s^{\min}_{i}}(\mathbf{p}_{-i})={\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}1\vee s^{\min}_{i}}(\Psi_{i}(p^{(Y_{n+i})}_{i},W_{i}))$
is a nonincreasing function of $p^{(Y_{n+i})}_{i}$. By Assumption 4 (ii) and applying Lemma H.1 (conditionally on $W_{i}$), we obtain
$\mathbb{E}\left[\frac{{\mathbf{1}\{p_{i}^{(Y_{n+i})}\leq\alpha({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}1\vee s^{\min}_{i}}(\mathbf{p}_{-i}))/m\}}}{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}1\vee s^{\min}_{i}}(\mathbf{p}_{-i})}\>\middle|\>W_{i}\right]\leq\frac{\alpha}{m},$
@@
C.3 Examining Assumption 4
We show here that the full-calibrated and class-calibrated $p$-value families satisfy Assumption 4 in the iid and conditional models, respectively. For this, it is interesting to relax Assumption 1 by assuming that the score functions can use the covariates of the calibration+test samples in an exchangeable way, as suggested in Marandon et al., (2024); Gazin et al., (2024). This is useful for instance when the learning sample and the calibration+test samples are not based on the same distribution, so that the scores may be improved by using transfer learning; we refer to Gazin et al., (2024) for more details on this.